OpenStudy (anonymous):

What is the function rule for the graph showing the ordered pairs , , , and ?

OpenStudy (anonymous):

incomplete

OpenStudy (anonymous):

Oh, sorry Here it is. What is the function rule for the graph showing the ordered pairs (2,0), (3,1), (4,2), and (5,3)? y = 2x y = x + 2 y= x/2 y = x - 2